A spring stretches to 2.S cm when 20 N is hanged to it. If a 40 N is hanged it has moved to 5 cm. The spring constant of the spring in N/m is
 
a. 700
b.800
c. 850
d.950 

Grade:12

1 Answers

mgayathri
6 years ago
i think it is 800n/m
m/k=l/g
equating basic formula of time period andsquaring the roots we get the above equation
on takingm=40N/10=4kg 
and l=5x10-2 and g=10.we get 800N/m